# Prompted Meta-Learning for Few-shot Knowledge Graph Completion

### v1

@@ 3 Problem Formulation
As defined above, few-shot KGC is a relation-specific task. The goal of a few-shot KGC model is to accurately predict new triplets associated with a relation $r$, when only a few triplets are provided for training. Thus, the model is trained based on the unit of tasks, where each task $\mathcal{T}_{r}$ is to predict new triplets associated with a specific relation $r$. The objective is to optimize the model based on a set of meta-training tasks $\mathcal{D}_{train}=\{\mathcal{T}_{i}\}_{i=1}^{I}$, where each meta-training task $\mathcal{T}_{r}$, corresponding to a specific relation $r$, consists of a support set $\mathcal{S}_{r}$ and a query set $\mathcal{Q}_{r}$, i.e. $\mathcal{T}_{r}=\{\mathcal{S}_{r},\mathcal{Q}_{r}\}$. During testing, the model is required to infer new triplets for unseen relations in the test set $\mathcal{D}_{test}=\{\mathcal{T}_{j}\}_{j=1}^{J}$, given a corresponding few-shot support set. The relations in the test set are unseen during training, i.e.$\mathcal{R}_{train}\cap\mathcal{R}_{test}=\varnothing$.
@@ 4.2 Learnable Meta-Semantic Prompt Pool for Knowledge Transfer
where $[\mathbf{h}_{si};\mathbf{t}_{si}]$ indicates the concatenated semantic embeddings for the head and tail entities for each support triplet $(h_{i},r,t_{i})$. $f_{\textrm{sa}}$ is a self-attention function that aggregates the embeddings of $K$ support triplets, assigning higher weights to more representative triplets in the task-semantic embedding.
@@ 4.4 Training Objective and Model Optimization
$\mathrm{score}(h_{i},t_{i})$ $=||\mathbf{h}_{pi}+\mathbf{mr}_{r}-\mathbf{t}_{pi}||_{2},$ (11)
where $||\mathbf{x}||_{2}$ represents the $\ell_{2}$ norm of vector $\mathbf{x}$, $\mathbf{h}_{i}$ and $\mathbf{t}_{i}$ are the head and tail entity embeddings generated by $\mathbf{h}_{i}=\mathbf{h}_{ri}+\mathbf{h}_{si}$, $\mathbf{t}_{i}=\mathbf{t}_{ri}+\mathbf{t}_{si}$. $\mathbf{h}_{pi}$ and $\mathbf{t}_{pi}$ are their corresponding projection vectors. $\mathbf{r}_{pi}$ is the projection vector of $\mathbf{mr}_{r}$, and $\mathbf{I}^{m\times n}$ is an identity matrix (Ji et al., 2015). Consider the entire support set, we can further define a loss function as follows:
@@
where $\gamma$ is a hyper-parameter that determines the margin to separate positive pairs from negative pairs. $\textrm{score}(h_{i},t^{\prime}_{i})$ calculates the score of a negative pair $(h_{i},t^{\prime}_{i})$ which results from negative sampling of the positive pair $(h_{i},t_{i})\in\mathcal{S}_{r}$, i.e.$(h_{i},r,t^{\prime}_{i})\notin\mathcal{G}$.
@@
$\mathrm{score}(h_{j},t_{j})=||\textbf{h}_{pj}+\mathbf{mr^{\prime}}_{r}-\textbf{t}_{pj}||_{2},$ (14)
@@ 5.3 Implementation Details
Table 4: Ablation on a) prompt tuning weight; b) the number of negative samples in contrastive loss; c) the size of MSP pool. $5$-shot setting on Nell-One are reported. Default settings are marked in gray.
@@ 5.6 Hyperparameter Analysis
We set $N$ from 00 to $1024$ in Table 1(b). PromptMeta performs the best when $N=1024$. As $N$ increases, the performance increases accordingly, suggesting different meta-semantics are becoming more effectively differentiated from one another.
We set the size of the MSP pool $M$ as 00, $32$, $64$, $128$ in Table 1(c). The best performance is achieved when $M=64$ for Nell-One. When $M$ increases to $128$, the significant performance drop suggests that the semantic information indeed exhibits shared patterns essential for knowledge transfer.

### v2

@@ 3 Problem Formulation
[…]
@@ 4.2 Learnable Meta-Semantic Prompt Pool for Knowledge Transfer
where $\mathbf{h}_{si};\mathbf{t}_{si}]$ indicates the concatenated semantic embeddings for the head and tail entities for each support triplet $(h_{i},r,t_{i})$. $f_{\textrm{sa}}$ is a self-attention function that aggregates the embeddings of $K$ support triplets, assigning higher weights to more representative triplets in the task-semantic embedding.
@@ 4.4 Training Objective and Model Optimization
[…]
where $\mathbf{x}||_{2}$ represents the $\ell_{2}$ norm of vector $\mathbf{x}$, $\mathbf{h}_{i}$ and $\mathbf{t}_{i}$ are the head and tail entity embeddings generated by $\mathbf{h}_{i}=\mathbf{h}_{ri}+\mathbf{h}_{si}$, $\mathbf{t}_{i}=\mathbf{t}_{ri}+\mathbf{t}_{si}$. $\mathbf{h}_{pi}$ and $\mathbf{t}_{pi}$ are their corresponding projection vectors. $\mathbf{r}_{pi}$ is the projection vector of $\mathbf{mr}_{r}$, and $\mathbf{I}^{m\times n}$ is an identity matrix (Ji et al., 2015). Consider the entire support set, we can further define a loss function as follows:
@@
[…]
@@
[…]
@@ 5.3 Implementation Details
Table 4: Ablation on a) prompt tuning weight; b) the number of negative samples in contrastive loss; c) the size of MSP pool. 55-shot setting on Nell-One are reported. Default settings are marked in gray.
@@ 5.6 Hyperparameter Analysis
We set $N$ from 0 to $1024$ in Table 1(b). PromptMeta performs the best when $N=1024$. As $N$ increases, the performance increases accordingly, suggesting different meta-semantics are becoming more effectively differentiated from one another.
We set the size of the MSP pool $M$ as 0, $32$, $64$, $128$ in Table 1(c). The best performance is achieved when $M=64$ for Nell-One. When $M$ increases to $128$, the significant performance drop suggests that the semantic information indeed exhibits shared patterns essential for knowledge transfer.